# Strong Normalization for the Safe Fragment of a Minimal Rewrite Calculus: A Triple-Lexicographic Proof and a Conjecture on Full-Termination Limits for Pure Recursive Calculi (PRC)

### v4

@@ 10 Formalization structure
All theorem statements currently asserted in the build are formally proven without unproven postulates. Confluence theorems in Meta/Newman_Safe.lean are stated with a global local-join parameter ($\mathsf{locAll}$). The artifact discharges this parameter via locAll_safe from Meta/Confluence_Safe.lean to obtain unconditional confluentSafe and unique normal forms for SafeStep. The formal development (Lean 4, $\sim$6,000 LOC; dual-licensed for academic and commercial use) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
@@ 11 Conclusion
This paper presents the KO7 operator-only kernel, proving strong normalization for a guarded safe subrelation and formalizing a Newman confluence engine. These results yield a certified normalizer; unique normal forms follow unconditionally for the safe fragment. An explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is established for the full relation, motivating the SafeStep fragment. The decidability result explains the clash between single-level Gödel encodings and global termination. The impossibility of additive bumps under duplication is formally verified.
@@ A.1 Lean kernel location
A.2 Auxiliary orientation sketch (DM/MPO)
This subsection records historical orientation notes used for boundary analysis. It is not the canonical SafeStep termination certificate in this paper, which is the computable stack in Meta/ComputableMeasure.lean. An auxiliary triple $\widetilde{\mu}^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,r_{\mathrm{aux}}(t))$ is ordered lexicographically:
@@ A.2 Auxiliary orientation sketch (DM/MPO)
  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
@@
It is emphasized that $\kappa^{M}$ aggregates via multiset operations (union $\cup$ and addition ++), not scalar numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and $r_{\mathrm{aux}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $r_{\mathrm{aux}}$ coordinate (right branch of Prod.Lex).
@@
MPO triple weight. $\mathrm{weight}(\mathrm{pair}\,a\,b)=(\mathrm{headRank}(a),\,\mathrm{size}(a),\,\mathrm{size}(b))$ with $\mathrm{headRank}(\mathrm{s}\,\_)=2$, else $1$. If $x$ is unit/pair then first components decrease (1¡2). If $x=\mathrm{s}\,t$ then tie on 2; the second components satisfy $\mathrm{size}(t)+1<\mathrm{size}(t)+2$.
@@
A.3 Newman scope (guarded safe relation)
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step has an explicit non-local-join witness at root, eqW void void; accordingly full-step confluence is not claimed.
@@ A.4 Module map
The verification logic is distributed across key modules for SN, normalizer, and confluence engine. Local confluence lemmas are proved separately. Note: context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment (SafeStepCtx) rather than the full relation.
@@ A.6 Local-join report (Confluence)
Appendix B Orientation Strategy
This summarizes the exact orientation strategy realized in the proofs:
  • All guarded rules are oriented in one computable lex order Lex3c over $(\delta\text{-flag},\kappa^{M},\tau)$.
  • The rule rec_succ is discharged by the outer $\delta\text{-flag}$ drop (1 $\to$ 0).
  • Multiset-sensitive branches (notably rec_zero) are discharged by DM drops on $\kappa^{M}$, while merge_cancel is discharged by $\tau$.
  • Tie branches are discharged by strict inequalities in the computable natural rank $\tau$.
  • Strong normalization follows from well-foundedness of Lex3c via inverse-image and yields wf_SafeStepRev_c.

### v5

@@ 10 Formalization structure
All theorem statements currently asserted in the build are formally proven without sorry, admit, or unsafe declarations. Confluence theorems in Meta/Newman_Safe.lean are stated with a global local-join parameter ($\mathsf{locAll}$). The artifact discharges this parameter via localJoin_all_safe from Meta/Confluence_Safe.lean, yielding locAll_safe in Meta/Newman_Safe.lean, to obtain unconditional confluentSafe and unique normal forms for SafeStep. The formal development (Lean 4, $\sim$6,000 LOC; dual-licensed for academic and commercial use) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
@@ 11 Conclusion
This paper presents the KO7 operator-only kernel, proving strong normalization for a guarded safe subrelation and formalizing a Newman confluence engine. These results yield a certified normalizer; unique normal forms follow unconditionally for the safe fragment. An explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is established for the full relation, motivating the SafeStep fragment. Impossibility results for twelve distinct strategy families (including additive, polynomial, weight-based, and precedence-based measures) are formally verified under duplication, and this evidence motivates a scoped conjectural boundary for unrestricted PRC termination under Definition 9.1.
@@ A.1 Lean kernel location
A.2 Layered Perspective (Conjectural Discussion)
This subsection records a discussion framework used to contextualize the PRC conjecture. It is not a proved reduction between formalisms.
Table 3: Conjectural layered comparison (discussion only).
Layer Calculus class Minimum requirement Classical barrier
0 (scoped conjecture) PRC Order + object-level recursor No known internally definable proof of unrestricted full termination
1 (Turing) General computation External tape + meta-level transition Halting undecidability
2 (Gödel) Arithmetic calculus Robinson Q (++ and $\times$) Incompleteness/consistency limits
Interpretation note
The table is a framing device: it compares the amount of external structure assumed in each setting. It does not claim a theorem stating that Layer 0 is formally reducible to or strictly below Layers 1 or 2.
A.3 Auxiliary orientation sketch (historical; external MPO comparator)
This subsection records historical orientation notes used for boundary analysis. It is not the canonical SafeStep termination certificate in this paper, which is the computable stack in Meta/ComputableMeasure.lean. Any MPO reference below is for external comparison only and is excluded from internally definable methods by Definition 9.1. An auxiliary triple $\widetilde{\mu}^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,r_{\mathrm{aux}}(t))$ is ordered lexicographically:
@@ A.3 Auxiliary orientation sketch (historical; external MPO comparator)
  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over an internal base numeric order on interpreted payloads, orienting redexes strictly above pieces; covers duplicators.
@@
It is emphasized that $\kappa^{M}$ aggregates via multiset operations (union $\cup$ and addition ++), not scalar numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component where applicable. For non-duplicating rules, $\kappa^{M}$ may tie by definitional equality and $r_{\mathrm{aux}}$ resolves the branch. In the canonical SafeStep certificate, merge_cancel is guarded by $\kappa^{M}(t)=0$, so strict decrease is discharged in the $r_{\mathrm{aux}}$$\tau$ coordinate. For unguarded eqW-refl analysis, a $\kappa$-branch split is used: when $\kappa^{M}(a)\neq 0$, a left-lex DM drop applies; when $\kappa^{M}(a)=0$, $\kappa^{M}$ ties by rfl and strict decrease is witnessed in the $r_{\mathrm{aux}}$ coordinate (right branch of Prod.Lex).
@@
Historical MPO comparator (external). $\mathrm{weight}(\mathrm{pair}\,a\,b)=(\mathrm{headRank}(a),\,\mathrm{size}(a),\,\mathrm{size}(b))$ with $\mathrm{headRank}(\mathrm{s}\,\_)=2$, else $1$. If $x$ is unit/pair then first components decrease (1¡2). If $x=\mathrm{s}\,t$ then tie on 2; the second components satisfy $\mathrm{size}(t)+1<\mathrm{size}(t)+2$.
@@ A.4 Module map
The verification logic is distributed across the following modules:
  • Kernel.lean — 7-constructor Trace inductive, 8-rule Step relation, reflexive-transitive closure Star.
  • Meta/SafeStep_Core.leanweight, kappaM, deltaFlag, guarded SafeStep relation.
  • Meta/ComputableMeasure.leantau, Lex3c, per-rule decrease lemmas (drop_R_*_c), wf_SafeStepRev_c.
  • Meta/Normalize_Safe.lean — Certified normalizer via WellFounded.fix; totality and soundness proofs.
  • Meta/SafeStep_Ctx.lean — Context closure SafeStepCtx for the safe fragment.
  • Meta/Confluence_Safe.lean — Local-join lemmas per root/context shape; localJoin_all_safe.
  • Meta/Newman_Safe.lean — Newman’s lemma instantiation; confluentSafe, unique normal forms.
  • Meta/Impossibility_Lemmas.lean — Machine-checked failed-measure catalog.
  • Meta/Conjecture_Boundary.lean — No-go theorems, GlobalOrients framework.
  • Meta/Operational_Incompleteness.lean
    Probe predicates (P1–P3), InternallyDefinableMeasure.
Context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment rather than the full relation.